light ray ′B′ travels extra distance 2L (for small angles)

has wavelength in coating λ_{coat} = (\frac{n_{1}}{n_{2}} ) λ_{air}

has phase shift of \frac{1}{2}λ at reflection with silicon if n_{2} < n_{3}

\frac{2L}{λ_{air}(\frac{n_{1}}{n_{2}} )} = N + \frac{1}{2} N = 0, 1, 2, 3, ...

L = \frac{1}{4} λ_{air} (\frac{n_{1}}{n_{2}} )(N − 1) N = 0, 1, 2, 3, ...